# Outlyingness Scores with Cluster Catch Digraphs

### v1

@@ 5 Summary and Conclusion
In this paper, we introduced two novel OSs based on CCDs: OOS and IOS. Both scores enhance the interpretability of outlier detection results. Both OSs employ graph-, density-, and distribution-based techniques, tailored to high-dimensional data with varying cluster shapes and intensities. The newly proposed OSs are able to identify both global and local outliers, and are invariant to data collinearity. The IOS method, in particular, exhibits robustness to the masking problem, which is a challenge for many outlier detection techniques. Through extensive Monte Carlo simulations and real-life data examples, we demonstrated the efficacy of both OSs, showcasing their superior performance in comparison to existing CCD-based methods and other state-of-the-art approaches.

### v2

@@ 5 Summary and Conclusion
In this paper, we introduced two novel OSs based on CCDs: OOS and IOS. Both scores enhance the interpretability of outlier detection results. Both OSs employ graph-, density-, and distribution-based techniques, tailored to high-dimensional data with varying cluster shapes and intensities. The newly proposed OSs can identify both global and local outliers and are invariant to data collinearity. The IOS method, in particular, exhibits robustness to the masking problem, which is a challenge for many outlier detection techniques. Through extensive Monte Carlo simulations and real-life data examples, we demonstrated the efficacy of both OSs, showcasing their superior performance in comparison to existing CCD-based methods and other state-of-the-art approaches.
@@ 6 Acknowledgements
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the author used ChatGPT (OpenAI) to assist with drafting, rephrasing, and refining portions of the text for clarity and conciseness. After using this tool, the author thoroughly reviewed and edited the content to ensure accuracy, coherence, and alignment with the scholarly intent of the work, and takes full responsibility for the content of the publication.